Super Carpeting Inc. just paid a dividend (D0D0) of $2.16, and its dividend is expected to grow at a constant rate (g) of 3.15% per year.

If the required return (rsrs) on Super’s stock is 7.88%, then the intrinsic, or theoretical market, value of Super’s shares is   per share.

Which of the following statements is true about the constant growth model?

When using a constant growth model to analyze a stock, if an increase in the required rate of return occurs while the growth rate remains the same, this will lead to an increased value of the stock.

When using a constant growth model to analyze a stock, if an increase in the required rate of return occurs while the growth rate remains the same, this will lead to a decreased value of the stock.

Use the constant growth model to calculate the appropriate values to complete the following statements about Super Carpeting Inc.:

If Super’s stock is in equilibrium, the current expected dividend yield on the stock will be   per share.
Super’s expected stock price one year from today will be   per share.
If Super’s stock is in equilibrium, the current expected capital gains yield on Super’s stock will be   per share.
Answer #1

When using a constant growth model to analyze a stock, if an increase in the required rate of return occurs while the growth rate remains the same, this will lead to a decreased value of the stock.

Current expected Dividend yield=r-g=7.88%-3.15%=4.73%
Expected stock price=D2/(r-g)=2.16*(1+3.15%)^2/(7.88%-3.15%)=48.58823
Current expected capital gains yield=g=3.15%